OpenStudy (h_hoff):

A bacterial culture isolated from sewage produced 64.1 mL of methane gas at 25oC and 7.80 x 102 torr. What would the volume of the methane be at STP?

OpenStudy (joannablackwelder):

P1V1/T1=P2V2/T2

OpenStudy (h_hoff):

ok

OpenStudy (h_hoff):

thanks

OpenStudy (joannablackwelder):

Make sure to use T in Kelvin. You're welcome. :-)

OpenStudy (h_hoff):

would you mind helping me with another question?

OpenStudy (joannablackwelder):

I can try.

OpenStudy (h_hoff):

A steel container with a volume that cannot change is filled with 125.0 atms of compressed air at 20.0 oC. It sits out in the sun all day, and its temperature rises to 65.0 oC. What is the new pressure?

OpenStudy (joannablackwelder):

P1/T1=P2/T2

OpenStudy (h_hoff):

thank you so much.

OpenStudy (joannablackwelder):

You're welcome. Again, change T into Kelvin.